In a recent year, the total sales at Restaurant A were $350,000 more then at Resteraunt B. Total sales at the two stores were 600,000,000 what were the sales for the year at each store?
Answer by rfer(16322)

A=x+350000
B=x
---------
x+x+350000=600000000
2x=599650000
x=299,825,000
x+350000=300,175,000